# Human as AI Mentor: Enhanced Human-in-the-loop Reinforcement Learning for Safe and Efficient Autonomous Driving

### v3

@@ 3.2.2 Implicit Intervention Mechanism
Assumption 3 (Model-based Traffic Flow Dynamics). We assume that the AV’s action $a_{\_}t^{AV}$ will persist for a short time interval $T$, and the following vehicles will adhere to the IDM (Treiber et al., 2000). This assumption is simplified and omits various factors present in real-world scenarios, such as multiple lanes, diverse driver behaviors, and varying vehicle parameters. Importantly, the traffic state we are predicting here is solely used for computing the disturbance cost and interpreting the impact of different agent actions on traffic flow.
@@
Remark 3: In practice, the current traffic flow state can be obtained through onboard sensors such as LIDAR and camera (Sheng et al., 2023). Additionally, it is possible to acquire velocity information for the following HVs of the leading AV at the current time via roadside units (RSUs) (Huang et al., 2023a; Lin et al., 2021; Olovsson et al., 2022). Furthermore, aside from model-based methods, we can also leverage deep learning techniques to enhance the prediction accuracy of future traffic flow states, as demonstrated in our previous works (Wu et al., 2020; Li et al., 2022a; Xu et al., 2023).
@@ 4.2.1 Vanilla Off-Policy Actor-Critic
Off-policy actor-critic is a well-known algorithm architecture in the RL field, which consists of two core components: actor and critic (Lillicrap et al., 2015). The actor’s task is to generate a policy; in other words, it determines the action to be taken in a given state. Conversely, the ‘critic’ is responsible for estimating the potential value of executing such an action in that state. This architecture is referred to as an ‘off-policy’, as its learning process can be independent of the current policy, learning instead from actions executed under different policies.
Under a policy $\pi$, the critic $Q$ is defined as follows (Lillicrap et al., 2015):
@@
The Eq. 13 assumes that the reward can reflect users’ intentions and effectively incentivize the desired behaviors of the agents. Reward functions are typically designed through trial-and-error methods. For instance,  Shi et al. (2021); Jiang et al. (2022); Wu et al. (2023) manually designed the reward function for autonomous driving, focusing on driving safety and comfort. On the other hand,  Huang et al. (2023b) employed IRL to infer the reward function from human demonstrations. However, as mentioned earlier, it is nearly impractical to encapsulate all driving behaviors facing various unforeseeable traffic scenarios in reward functions.
@@ 4.2.2 Removing the Reward Function
In a recent survey (Booth et al., 2023), it was found that 92% of surveyed expert-level RL practitioners reported using manual trial-and-error reward designs, with 89% indicating that their designed reward was suboptimal. Knox et al. (2023) developed eight sanity checks to identify flaws in reward functions and surveyed papers on RL-based AV published in top-tier venues, finding near-universal flaws in reward design for AVs. Thus, designing a reward function that is universally applicable to all scenarios for AVs is a complex and challenging task. Recent work has demonstrated that agents can learn complex goals by interacting with humans without a fixed reward function (Li et al., 2022c; Ibarz et al., 2018; Peng et al., 2023).
The distinction of this work from previous efforts is the removal of the reward function. Revisiting the primary goal in the driving policy learning task, we find that reward functions are not necessary, as our true aim is to implement human preference in learning behavior. We observe that when a human expert intervenes by taking over the agent, a clear signal is generated, indicating dissatisfaction with the agent’s current performance, whether due to unsafe actions or poorly performing behaviors. Conversely, if the human experts do not intervene, then the current states and the agent’s actions are deemed to be human-compatible (Peng et al., 2023). Based on this observation, we can use a proxy value function to denote human preference instead of the traditional reward function; this does not include reward but only calculates a proxy $Q$ values (Peng et al., 2023).
@@ 4.3.1 Learning from Explicit Intervention
As we can only learn from partial demonstrations in $\mathcal{B}$, this leads to a distribution shift problem. To address this issue, we adopt the recent CQL (Kumar et al., 2020) to train using off-policy data generated by humans. In detail, we sample data $\left(s_{\_}t,a_{\_}t^{AV},a_{\_}t^{human},I\left(s_{\_}t,a_{\_}t^{AV}\right)\right)$ from the replay buffer $\mathcal{B}$, and label the proxy $Q$ value of human action $a_{\_}t^{AV}$ with $\hat{Q}\left(s_{\_}t,a^{human}\right)$ and the agent action $a_{\_}t^{AV}$ with $\hat{Q}\left(s_{\_}t,a^{AV}\right)$.
@@ 4.3.2 Learning from Exploration with Entropy Regularization
When the learning agent fails to adequately explore the human-preferred subspace during free exploratory sampling, states that trigger high proxy values are infrequently encountered. This poses a challenge in propagating the proxy value back to previous states, impeding the learning process. To encourage exploration, we utilize the entropy regularization technique (Haarnoja et al., 2018), creating an auxiliary signal for updating the proxy value function:
@@
Different from previous offline RL works (Nair et al., 2018; Kumar et al., 2020), which train from a fixed dataset without access to closed-loop feedback, HAIM-DRL utilizes online exploration and partial human demonstration data. Furthermore, HAIM-DRL ensures continuity in state visitation between the human expert and the agent, effectively mitigating the potential OOD issues mentioned earlier in limitation (b).
@@ 4.3.3 Learning from Implicit Intervention
Eq. 2 illustrates that even minor changes in the state of the leading AV can lead to disturbances in the following HVs, disrupting the overall traffic flow balance. In the transportation community, there are many indicators used to assess the performance of the traffic system, such as acceleration, headway, average velocity, and time to collision (TTC) (Shi et al., 2023; Jiang et al., 2022; Zhu et al., 2020; Mahdinia et al., 2021). In particular, Sharma et al. (2021) proposed using a wavelet energy-based method to calculate traffic oscillation metrics (i.e., oscillation duration and amplitude) to analyze traffic flow disturbance. The main idea of their work is to identify significant changes in velocity. Inspired by this, and building upon Eq. 3, we propose a measurement method for disturbance cost, which can quantify the potential downstream impact of AVs on the following vehicles.
@@ 5             $v_{\_}{t+T,i}^{HV}\leftarrow$ Compute the desire HVs’ velocity based on Eq 9;
6      $v_{\_}{t+T,N}^{HV}\leftarrow\frac{1}{NT}\sum_{\_}{i=1}^{N}v_{\_}{t+T,i}^{HV}$ Compute the velocity deviation: $\Delta v\leftarrow v_{\_}{t+T,N}^{HV}-v_{\_}{t,N}^{HV}$ if $acc_{\_}t^{AV}\geq\lambda\text{ and }acc_{\_}{t-1}^{AV}<\lambda$ then
@@ 4.3.4 Reducing the Human Mentor’s Cognitive Load
The learning agent faces the risk of excessive dependence on human if the frequency of human takeover is not restricted, leading to failure when evaluating the agent without human supervision (Peng et al., 2022; Li et al., 2022c; Wu et al., 2023). This risk arises because $\hat{Q}(s_{\_}t,a_{\_}t)$ represents the proxy $Q$ value of the mixed behavior policy $\pi_{\_}{mix}$, rather than the AV’s learning policy $\pi_{\_}{AV}$, owing to the presence of human takeover. In this case, there’s a likelihood that the agent may select actions that go against human wishes, such as veering off the road when approaching the boundary. This behavior compels human to intervene and offer corrective demonstrations. In consequence, the AV’s degree of automation remains low, leading to the human mentor experiencing fatigue due to the continual need to provide demonstrations.
Definition 7 (Takeover Cost). To mitigate human mentor’s cognitive load and enhance the AV’s level of automation, we apply a mild penalty to agent behaviors that trigger human explicit intervention. To achieve this, we employ the cosine similarity between the actions of the agent and those of a human as the takeover cost, which is formulated as follows (Li et al., 2022c):
@@
The agent incurs a significant penalty only when there’s a substantial difference between the cosine similarity of its actions and those of humans. Li et al. (2022c) illustrated that employing this cosine similarity yields superior results compared to a fixed cost, such as using a ‘+1’ penalty. Additionally, we attribute the takeover cost to the agent only during the initial step of human takeover. This approach is grounded in the observation that human takeover, triggered by a specific action $a^{AV}_{\_}t$, signify a deviation from human intentions at that particular moment.
@@
3             $a_{\_}t^{AV}\sim\pi_{\_}{AV}(\cdot|s_{\_}t)$ // Retrieve agent’s action Execute $a_{\_}t^{AV}$ and observe new state $s_{\_}{t+1}$ $I(s_{\_}t,a_{\_}t^{AV})\leftarrow$ Human expert determines whether to intervene by observing current state $s_{\_}t$ if $I(s_{\_}t,a_{\_}t^{AV})$ is True then
4                   $a_{\_}t^{human}\leftarrow\pi_{\_}{human}(\cdot|s_{\_}t)$ // Retrieve human’s action Execute $a_{\_}t^{human}$ within the environment
@@ 5.1.1 Experimental Scenarios
Due to the potential risks associated with involving human experts in physical experiments, we conducted benchmarking using a driving simulator. Specifically, we employed the latest lightweight driving simulator, MetaDrive (Li et al., 2022b), which enables the evaluation of safety and generalizability in unknown environments. MetaDrive is implemented using Panda3D and the Bullet Engine, providing efficient and accurate physics-based 3D dynamics. Some generated driving scenarios are depicted in Fig. 5 (a). While the MetaDrive setup is primarily described in this section, experiments were also conducted using the CARLA simulator (Dosovitskiy et al., 2017) to verify the transferability of the proposed method. For interaction with the CARLA core, we utilize the CARLA client wrapper implemented in DI-drive (drive Contributors, 2021), as depicted in Fig. 5 (b).
To sufficiently encompass dangerous and long-tail scenarios, we used MetaDrive to generate a diverse array of driving scenarios, each training session comprising 50 different scenarios. As depicted in Fig. 5 (a), this encompasses road scenarios made up of various typical block types such as straight, ramp, fork, roundabout, curve, T-intersection, and intersection. Surrounding vehicles are controlled using two distinct models: the IDM for longitudinal dynamics, and the minimizing overall braking induced by lane change model (MOBIL) (Kesting et al., 2007) for lateral movement. Furthermore, we introduced traffic flows at varying densities to interact with the AV. To enhance the complexity of the driving task, each scene included randomly placed obstacles, such as moving traffic vehicles, stationary traffic cones, and triangular warning signs.
@@ 5.1.3 Observation, Reward, Cost, and Generalization Ability
Environmental Cost. Each collision with traffic vehicles or obstacles results in an environmental cost of ‘+1’. Note that HAIM-DRL does not have access to this cost during training. This cost is utilized for training safe RL methods and for testing the safety of trained policies. Detailed definitions of reward and costs can be found in Li et al. (2022b). The episodic cost, referred to as ‘Safety Violation’, serves as an indicator of a policy’s safety level.
Generalization Ability. Success rate is a more appropriate criterion for assessing generalization compared to episodic reward. This is because the evaluated scenes exhibit diverse attributes, such as road length and traffic density, which result in significant reward fluctuations across different scenes (Li et al., 2022b).
@@ 5.1.4 Implementation Details
As emphasized in Assumption 1, the HAIM-DRL framework requires only a single experienced human expert for implementation. In this experiment, as shown in Figs. 5 (a), a human participant with both United States and Chinese driving licenses and 7 years of driving experience was invited to supervise the learning agent’s real-time exploration with hands on the steering wheel. Upon perceiving an impending dangerous situation, the human takes over by pressing the paddle beside the wheel, thereby gaining control of the vehicle through steering and pedal manipulation. Additionally, we assume that there are $N$ HVs following the AV agent, driving according to the IDM. Too large an ‘$N$’ can complicate traffic management, whereas too small an ‘$N$’ might oversimplify real-world traffic dynamics. Similarly, a very high ‘$X$’ can overwhelm the AV with unpredictable elements, while a low ‘$X$’ may not adequately test its adaptability. In our study, the ‘$N$’ value aligns with Shi et al. (2023), and ‘$X$’ with Li et al. (2022c). The B describes the detailed hyperparameter settings used in this experiment.
@@
RL SAC-RS (Haarnoja et al., 2018) 2.78K $\pm$ 0.97K 1M 384.56 $\pm$37.5 0.87 $\pm$1.47 0.015 $\pm$ 0.005 0.83 $\pm$0.32
PPO-RS (Schulman et al., 2017) 27.51K $\pm$3.86K 1M 305.41 $\pm$14.23 4.12 $\pm$1.24 0.021 $\pm$ 0.014 0.67$\pm$0.12
Safe RL SAC-Lag (Ha et al., 2021) 1.98K $\pm$ 0.75K 1M 352.46 $\pm$108.78 0.78 $\pm$0.58 0.019 $\pm$ 0.007 0.71 $\pm$0.79
PPO-Lag (Stooke et al., 2020) 15.46K $\pm$ 5.13K 1M 298.98 $\pm$50.99 3.28 $\pm$0.38 0.025 $\pm$ 0.016 0.52 $\pm$0.27
CPO (Achiam et al., 2017) 4.36K $\pm$2.22K 1M 194.06 $\pm$108.86 1.71 $\pm$1.02 - 0.21 $\pm$0.29
Offline RL CQL (Kumar et al., 2020) - 49K 116.45 $\pm$34.94 3.68 $\pm$7.61 0.007 $\pm$ 0.0006 0.13 $\pm$0.09
IL BC (Bain and Sammut, 1995) - 49K 36.13 $\pm$10.66 1.05 $\pm$0.54 0.012 $\pm$ 0.0170 0.01 $\pm$0.02
GAIL (Ho and Ermon, 2016) 3.68K $\pm$3.17K 49K 108.36 $\pm$16.08 4.18 $\pm$1.25 0.001 $\pm$ 0.0009 0.03 $\pm$0.01
Conventional Human-in-the-loop HG-DAgger (Kelly et al., 2019) 35.58 50K 106.21 2.63 0.108 0.04
IWR (Mandlekar et al., 2020) 69.74 50K 298.87 3.61 0.122 0.61
Human-AI Copilot HACO (Li et al., 2022c) 30.05 $\pm$10.89 30K 350.01 $\pm$9.72 0.78 $\pm$0.85 0.038 $\pm$ 0.0083 0.83 $\pm$0.07

### v5

@@ 3.2.2 Implicit Intervention Mechanism
Assumption 3 (Model-based Traffic Flow Dynamics). We assume that the AV’s action $a_{\_}t^{AV}$ will persist for a short time interval $T$, and the following vehicles will adhere to the IDM [Treiber et al., 2000]. This assumption is simplified and omits various factors present in real-world scenarios, such as multiple lanes, diverse driver behaviors, and varying vehicle parameters. Importantly, the traffic state we are predicting here is solely used for computing the disturbance cost and interpreting the impact of different agent actions on traffic flow.
@@
Remark 3: In practice, the current traffic flow state can be obtained through onboard sensors such as LIDAR and camera [Sheng et al., 2023]. Additionally, it is possible to acquire velocity information for the following HVs of the leading AV at the current time via roadside units (RSUs) [Huang et al., 2023a, Lin et al., 2021, Olovsson et al., 2022]. Furthermore, aside from model-based methods, we can also leverage deep learning techniques to enhance the prediction accuracy of future traffic flow states, as demonstrated in our previous works [Wu et al., 2020, Li et al., 2022a, Xu et al., 2023].
@@ 4.2.1 Vanilla Off-Policy Actor-Critic
Off-policy actor-critic is a well-known algorithm architecture in the RL field, which consists of two core components: actor and critic [Lillicrap et al., 2015]. The actor’s task is to generate a policy; in other words, it determines the action to be taken in a given state. Conversely, the ‘critic’ is responsible for estimating the potential value of executing such an action in that state. This architecture is referred to as an ‘off-policy’, as its learning process can be independent of the current policy, learning instead from actions executed under different policies.
Under a policy $\pi$, the critic $Q$ is defined as follows [Lillicrap et al., 2015]:
@@
The Eq. 13 assumes that the reward can reflect users’ intentions and effectively incentivize the desired behaviors of the agents. Reward functions are typically designed through trial-and-error methods. For instance,  Shi et al. [2021], Jiang et al. [2022], Wu et al. [2023] manually designed the reward function for autonomous driving, focusing on driving safety and comfort. On the other hand,  Huang et al. [2023b] employed IRL to infer the reward function from human demonstrations. However, as mentioned earlier, it is nearly impractical to encapsulate all driving behaviors facing various unforeseeable traffic scenarios in reward functions.
@@ 4.2.2 Removing the Reward Function
In a recent survey [Booth et al., 2023], it was found that 92% of surveyed expert-level RL practitioners reported using manual trial-and-error reward designs, with 89% indicating that their designed reward was suboptimal. Knox et al. [2023] developed eight sanity checks to identify flaws in reward functions and surveyed papers on RL-based AV published in top-tier venues, finding near-universal flaws in reward design for AVs. Thus, designing a reward function that is universally applicable to all scenarios for AVs is a complex and challenging task. Recent work has demonstrated that agents can learn complex goals by interacting with humans without a fixed reward function [Li et al., 2022c, Ibarz et al., 2018, Peng et al., 2023].
The distinction of this work from previous efforts is the removal of the reward function. Revisiting the primary goal in the driving policy learning task, we find that reward functions are not necessary, as our true aim is to implement human preference in learning behavior. We observe that when a human expert intervenes by taking over the agent, a clear signal is generated, indicating dissatisfaction with the agent’s current performance, whether due to unsafe actions or poorly performing behaviors. Conversely, if the human experts do not intervene, then the current states and the agent’s actions are deemed to be human-compatible [Peng et al., 2023]. Based on this observation, we can use a proxy value function to denote human preference instead of the traditional reward function; this does not include reward but only calculates a proxy $Q$ values [Peng et al., 2023].
@@ 4.3.1 Learning from Explicit Intervention
As we can only learn from partial demonstrations in $\mathcal{B}$, this leads to a distribution shift problem. To address this issue, we adopt the recent CQL [Kumar et al., 2020] to train using off-policy data generated by humans. In detail, we sample data $\left(s_{\_}t,a_{\_}t^{AV},a_{\_}t^{human},I\left(s_{\_}t,a_{\_}t^{AV}\right)\right)$ from the replay buffer $\mathcal{B}$, and label the proxy $Q$ value of human action $a_{\_}t^{AV}$ with $\hat{Q}\left(s_{\_}t,a^{human}\right)$ and the agent action $a_{\_}t^{AV}$ with $\hat{Q}\left(s_{\_}t,a^{AV}\right)$.
@@ 4.3.2 Learning from Exploration with Entropy Regularization
When the learning agent fails to adequately explore the human-preferred subspace during free exploratory sampling, states that trigger high proxy values are infrequently encountered. This poses a challenge in propagating the proxy value back to previous states, impeding the learning process. To encourage exploration, we utilize the entropy regularization technique [Haarnoja et al., 2018], creating an auxiliary signal for updating the proxy value function:
@@
Different from previous offline RL works [Nair et al., 2018, Kumar et al., 2020], which train from a fixed dataset without access to closed-loop feedback, HAIM-DRL utilizes online exploration and partial human demonstration data. Furthermore, HAIM-DRL ensures continuity in state visitation between the human expert and the agent, effectively mitigating the potential OOD issues mentioned earlier in limitation (b).
@@ 4.3.3 Learning from Implicit Intervention
Eq. 2 illustrates that even minor changes in the state of the leading AV can lead to disturbances in the following HVs, disrupting the overall traffic flow balance. In the transportation community, there are many indicators used to assess the performance of the traffic system, such as acceleration, headway, average velocity, and time to collision (TTC) [Shi et al., 2023, Jiang et al., 2022, Zhu et al., 2020, Mahdinia et al., 2021]. In particular, Sharma et al. [2021] proposed using a wavelet energy-based method to calculate traffic oscillation metrics (i.e., oscillation duration and amplitude) to analyze traffic flow disturbance. The main idea of their work is to identify significant changes in velocity. Inspired by this, and building upon Eq. 3, we propose a measurement method for disturbance cost, which can quantify the potential downstream impact of AVs on the following vehicles.
@@ 8             $v_{\_}{t+T,i}^{HV}\leftarrow$ Compute the desire HVs’ velocity based on Eq 9;
9      $v_{\_}{t+T,N}^{HV}\leftarrow\frac{1}{NT}\sum_{\_}{i=1}^{N}v_{\_}{t+T,i}^{HV}$
10
11      Compute the velocity deviation: $\Delta v\leftarrow v_{\_}{t+T,N}^{HV}-v_{\_}{t,N}^{HV}$
12
13      if $acc_{\_}t^{AV}\geq\lambda\text{ and }acc_{\_}{t-1}^{AV}<\lambda$ then
@@ 4.3.4 Reducing the Human Mentor’s Cognitive Load
The learning agent faces the risk of excessive dependence on human if the frequency of human takeover is not restricted, leading to failure when evaluating the agent without human supervision [Peng et al., 2022, Li et al., 2022c, Wu et al., 2023]. This risk arises because $\hat{Q}(s_{\_}t,a_{\_}t)$ represents the proxy $Q$ value of the mixed behavior policy $\pi_{\_}{mix}$, rather than the AV’s learning policy $\pi_{\_}{AV}$, owing to the presence of human takeover. In this case, there’s a likelihood that the agent may select actions that go against human wishes, such as veering off the road when approaching the boundary. This behavior compels human to intervene and offer corrective demonstrations. In consequence, the AV’s degree of automation remains low, leading to the human mentor experiencing fatigue due to the continual need to provide demonstrations.
Definition 7 (Takeover Cost). To mitigate human mentor’s cognitive load and enhance the AV’s level of automation, we apply a mild penalty to agent behaviors that trigger human explicit intervention. To achieve this, we employ the cosine similarity between the actions of the agent and those of a human as the takeover cost, which is formulated as follows [Li et al., 2022c]:
@@
The agent incurs a significant penalty only when there’s a substantial difference between the cosine similarity of its actions and those of humans. Li et al. [2022c] illustrated that employing this cosine similarity yields superior results compared to a fixed cost, such as using a ‘+1’ penalty. Additionally, we attribute the takeover cost to the agent only during the initial step of human takeover. This approach is grounded in the observation that human takeover, triggered by a specific action $a^{AV}_{\_}t$, signify a deviation from human intentions at that particular moment.
@@
3             $a_{\_}t^{AV}\sim\pi_{\_}{AV}(\cdot|s_{\_}t)$ // Retrieve agent’s action
4             Execute $a_{\_}t^{AV}$ and observe new state $s_{\_}{t+1}$
5             $I(s_{\_}t,a_{\_}t^{AV})\leftarrow$ Human expert determines whether to intervene by observing current state $s_{\_}t$
6             if $I(s_{\_}t,a_{\_}t^{AV})$ is True then
7                   $a_{\_}t^{human}\leftarrow\pi_{\_}{human}(\cdot|s_{\_}t)$ // Retrieve human’s action
8                   Execute $a_{\_}t^{human}$ within the environment
9
@@ 5.1.1 Experimental Scenarios
Due to the potential risks associated with involving human experts in physical experiments, we conducted benchmarking using a driving simulator. Specifically, we employed the latest lightweight driving simulator, MetaDrive [Li et al., 2022b], which enables the evaluation of safety and generalizability in unknown environments. MetaDrive is implemented using Panda3D and the Bullet Engine, providing efficient and accurate physics-based 3D dynamics. Some generated driving scenarios are depicted in Fig. 5 (a). While the MetaDrive setup is primarily described in this section, experiments were also conducted using the CARLA simulator [Dosovitskiy et al., 2017] to verify the transferability of the proposed method. For interaction with the CARLA core, we utilize the CARLA client wrapper implemented in DI-drive [drive Contributors, 2021], as depicted in Fig. 5 (b).
To sufficiently encompass dangerous and long-tail scenarios, we used MetaDrive to generate a diverse array of driving scenarios, each training session comprising 50 different scenarios. As depicted in Fig. 5 (a), this encompasses road scenarios made up of various typical block types such as straight, ramp, fork, roundabout, curve, T-intersection, and intersection. Surrounding vehicles are controlled using two distinct models: the IDM for longitudinal dynamics, and the minimizing overall braking induced by lane change model (MOBIL) [Kesting et al., 2007] for lateral movement. Furthermore, we introduced traffic flows at varying densities to interact with the AV. To enhance the complexity of the driving task, each scene included randomly placed obstacles, such as moving traffic vehicles, stationary traffic cones, and triangular warning signs.
@@ 5.1.3 Observation, Reward, Cost, and Generalization Ability
Environmental Cost. Each collision with traffic vehicles or obstacles results in an environmental cost of ‘+1’. Note that HAIM-DRL does not have access to this cost during training. This cost is utilized for training safe RL methods and for testing the safety of trained policies. Detailed definitions of reward and costs can be found in Li et al. [2022b]. The episodic cost, referred to as ‘Safety Violation’, serves as an indicator of a policy’s safety level.
Generalization Ability. Success rate is a more appropriate criterion for assessing generalization compared to episodic reward. This is because the evaluated scenes exhibit diverse attributes, such as road length and traffic density, which result in significant reward fluctuations across different scenes [Li et al., 2022b].
@@ 5.1.4 Implementation Details
As emphasized in Assumption 1, the HAIM-DRL framework requires only a single experienced human expert for implementation. In this experiment, as shown in Figs. 5 (a), a human participant with both United States and Chinese driving licenses and 7 years of driving experience was invited to supervise the learning agent’s real-time exploration with hands on the steering wheel. Upon perceiving an impending dangerous situation, the human takes over by pressing the paddle beside the wheel, thereby gaining control of the vehicle through steering and pedal manipulation. Additionally, we assume that there are $N$ HVs following the AV agent, driving according to the IDM. Too large an ‘$N$’ can complicate traffic management, whereas too small an ‘$N$’ might oversimplify real-world traffic dynamics. Similarly, a very high ‘$X$’ can overwhelm the AV with unpredictable elements, while a low ‘$X$’ may not adequately test its adaptability. In our study, the ‘$N$’ value aligns with Shi et al. [2023], and ‘$X$’ with Li et al. [2022c]. The B describes the detailed hyperparameter settings used in this experiment.
@@
RL SAC-RS [Haarnoja et al., 2018] 2.78K $\pm$ 0.97K 1M 384.56 $\pm$37.5 0.87 $\pm$1.47 0.015 $\pm$ 0.005 0.83 $\pm$0.32
PPO-RS [Schulman et al., 2017] 27.51K $\pm$3.86K 1M 305.41 $\pm$14.23 4.12 $\pm$1.24 0.021 $\pm$ 0.014 0.67$\pm$0.12
Safe RL SAC-Lag [Ha et al., 2021] 1.98K $\pm$ 0.75K 1M 352.46 $\pm$108.78 0.78 $\pm$0.58 0.019 $\pm$ 0.007 0.71 $\pm$0.79
PPO-Lag [Stooke et al., 2020] 15.46K $\pm$ 5.13K 1M 298.98 $\pm$50.99 3.28 $\pm$0.38 0.025 $\pm$ 0.016 0.52 $\pm$0.27
CPO [Achiam et al., 2017] 4.36K $\pm$2.22K 1M 194.06 $\pm$108.86 1.71 $\pm$1.02 - 0.21 $\pm$0.29
Offline RL CQL [Kumar et al., 2020] - 49K 116.45 $\pm$34.94 3.68 $\pm$7.61 0.007 $\pm$ 0.0006 0.13 $\pm$0.09
IL BC [Bain and Sammut, 1995] - 49K 36.13 $\pm$10.66 1.05 $\pm$0.54 0.012 $\pm$ 0.0170 0.01 $\pm$0.02
GAIL [Ho and Ermon, 2016] 3.68K $\pm$3.17K 49K 108.36 $\pm$16.08 4.18 $\pm$1.25 0.001 $\pm$ 0.0009 0.03 $\pm$0.01
Conventional Human-in-the-loop HG-DAgger [Kelly et al., 2019] 35.58 50K 106.21 2.63 0.108 0.04
IWR [Mandlekar et al., 2020] 69.74 50K 298.87 3.61 0.122 0.61
Human-AI Copilot HACO [Li et al., 2022c] 30.05 $\pm$10.89 30K 350.01 $\pm$9.72 0.78 $\pm$0.85 0.038 $\pm$ 0.0083 0.83 $\pm$0.07